If a cube of a certain metal increases its volume from 15 cm3 to 17 cm3 as the temperature increases from 12°C to 28°C, what is the metal's linear expansion coefficient?
Answer #1

linear expansion coef = change in volume/ 3 * temp change * volume

= (17 - 15) / ( 3 * 28-12 * 15

= 2.78 * 10^-3
Answer #2

Del V= 2 cm3

Vo=15

del T= 28-12= 16

coefficient= Del V / Vo * del T

= 2/15 * 16

= 0.0083333